OpenStudy (anonymous):

How do I get to step 3? 2(a-4)+15=13 Step 1 2a-8+15=13 Step 2

OpenStudy (bahrom7893):

2a + 7 = 13 2a = 13 - 7 2a = 6 a = 6/2 a = 3